a mother is 60 years old and her child is 24 years old. when in the past was the product of their age 160?
Answer by stanbon(75887)

a mother is 60 years old and her child is 24 years old. when in the past was the product of their age 160?
----
Equation:
(60-x)(24-x) = 160
x^2-84x+1440 = 160
---
x^2-84x+1280 = 0
x = 20 or x = 64
---
Acceptable answer:
x = 20 years
======================
Cheers,
Stan H.
======================